Bunnell Corporation is a manufacturer that uses job-order costing. On January 1, the companys inventory balances were as folJournal entry worksheet Record the manufactured goods completed during this year. Note: Enter debits before credits. Transact7. What is the ending balance in Work in Process? Work in Process Beg. Bal. End. Bal. 0 08. What is the total amount of actual manufacturing overhead cost incurred during the year? Total actual manufacturing overhe9. Is manufacturing overhead underapplied or overapplied for the year? By how much?
Answer #1

Journal entry

No General Journal Debit credit
Finished goods 1372600
Work in process 1372600
(To record goods completed)

7) Ending work in process

Beg WIP 22200 Finished goods 1372600
Direct material used 569400
Direct labor 410000
Manufacturing overhead 471500
End WIP 100500

8) Actual manufacturing overhead = 150000+310000 = 460000

9) Over applied overhead = 471500-460000 = 11500